-2x+3y=14
x-4y=-12

Answer by mananth(16946)

-2x+3y=14.............1
x-4y=-12 .............2
Eliminate y
multiply (1)by 4
Multiply (2) by 3
-8x+12y=56
3x-12y=-36
Add the two equations
-5x = 20
/-5
x=-4.00
plug value of x in (1)
-2x+3y= 14
8+3y=14
3y=14-8
3y=6
y=2